OpenStudy (anonymous):

[8.06] Identify the factors of 9x2 + 24x + 16. (3x + 4)(3x + 4) (3x − 4)(3x − 4) (3x + 2)(3x + 8) (3x − 8)(3x − 2)

OpenStudy (anonymous):

another perfect square,

OpenStudy (anonymous):

since \(16=4^2\), \(9x^2=(3x)^2\) and \(2\times 4\times 3=24\) it is a perfect square, namely \((3x+4)^2\)

OpenStudy (anonymous):

oh, and it is NOT C this time wow

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Is it A? right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

yw

OpenStudy (cwrw238):

to expand these you multiply the 2nd bracket by the first term in first bracket then multiply 2nd by second term in first bracket: (3x + 4)(3x + 4) = 3x(3x + 4) + 4(3x + 4) = 9x^2 + 12x + 12x + 16 = 9x^2 + 24x + 16

OpenStudy (anonymous):

you have virtual school

OpenStudy (anonymous):

yes and it suck

OpenStudy (anonymous):

same here todays my last day to finish everything for math and it sucks so bad what sate are u doing for @Felicia123

OpenStudy (anonymous):

Deltona High School

OpenStudy (anonymous):

cool im doing it for florida high but im in middle school